# Analog Computing for Signal Processing and Communications – Part II: Toward Gigantic MIMO Beamforming

### v2

@@ III-C Computing the Special Cases of the LMMSE Estimator with a MiLAC
The expressions of the five special cases of the LMMSE estimator can be efficiently returned on the output vector $\mathbf{v}_{2}$ of a MiLAC having $N=Y$ input ports and $M=X$ output ports with no input, i.e., $P=X+Y$ total ports. This is achieved by exploiting (4) and (6), as discussed for the LMMSE estimator in Part I of this paper. The input vector of the MiLAC is set as $\mathbf{u}=\mathbf{y}$ and the tunable admittance components $\{Y_{i,k}\}_{i,k=1}^{P}$ are set according to (7) such that the blocks of the matrix $\mathbf{P}$ give the desired expression in $\mathbf{v}_{2}$. In detail, the expressions of the special cases in (10)-(17) can be computed by setting the matrix $\mathbf{P}$ as illustrated in Tabs. II and III, depending on the invertibility of the blocks $\mathbf{P}_{11}$ and $\mathbf{P}_{22}$. When the block $\mathbf{P}_{11}$ is invertible, $\mathbf{v}_{2}$ is given by (4), as reported in Tab. II, while when the block $\mathbf{P}_{22}$ is invertible, $\mathbf{v}_{2}$ is given by (6), as reported in Tab. III. Note that GLS and OLS are computed via a matrix $\mathbf{P}$ whose blocks $\mathbf{P}_{11}$ and $\mathbf{P}_{22}$ are not both invertible. Thus, for GLS and OLS the expressions in Tabs. II and III are not equivalent in general, while they are equivalent for the LMMSE estimator and the other special cases.
@@
The five special cases of the LMMSE estimator can be efficiently computed using a MiLAC, which offers significantly reduced computational complexity compared to a digital computer. This is because a MiLAC directly performs matrix-matrix multiplications and matrix inversion operations in the analog domain. On the one hand, the computational complexity of computing the special cases with a MiLAC is driven by the complexity of computing (7) for a given matrix $\mathbf{P}$. Thus, the complexity required to compute GLS, RLS, and OLS with a MiLAC is the same as the complexity of computing the LMMSE estimator, i.e., $6XY$ real operations, as discussed in Part I of this paper. Besides, computing (7) for the GMF and OMF require approximately $2XY$ real operations for both cases $i\neq k$ and $i=k$, resulting in a total of $4XY$ real operations, where we assumed that the operations involving $\mathbf{C}_{\mathbf{x}}$ and $\mathbf{C}_{\mathbf{n}}$ are precomputed offline. On the other hand, a digital computer calculates the GLS with the same complexity as the LMMSE estimator due to the two matrix-matrix products and matrix inversion, i.e., $8(XY^{2}+X^{2}Y+\min\{X^{3},Y^{3}\}/3)$, as discussed in Part I of this paper. The RLS and OLS can be computed with reduced complexity as they require only one matrix-matrix product and a matrix inversion, namely $8\min\{X^{2}Y+X^{3}/3,XY^{2}+Y^{3}/3\}$. Besides, the GMF and OMF only require three and one matrix-vector products, respectively, resulting in a complexity of $8(X^{2}+XY+Y^{2})$ and $8XY$ real operations, respectively.
@@
LMMSE $\mathbf{C}_{\mathbf{x}}\mathbf{H}^{H}\left(\mathbf{H}\mathbf{C}_{\mathbf{x}}\mathbf{H}^{H}+\mathbf{C}_{\mathbf{n}}\right)^{-1}\mathbf{u}$ [±$\begin{bmatrix}\pm\mathbf{C}_{\mathbf{n}}&\mathbf{H}\\ \mathbf{H}^{H}&\mp\mathbf{C}_{\mathbf{x}}^{-1}\end{bmatrix}$
@@
GMF $\mathbf{C}_{\mathbf{x}}\mathbf{H}^{H}\mathbf{C}_{\mathbf{n}}^{-1}\mathbf{u}$ [±$\begin{bmatrix}\pm\mathbf{C}_{\mathbf{n}}&\mathbf{0}_{Y\times X}\\ \mathbf{H}^{H}&\mp\mathbf{C}_{\mathbf{x}}^{-1}\end{bmatrix}$
RLS $\mathbf{H}^{H}\left(\mathbf{H}\mathbf{H}^{H}+\lambda\mathbf{I}_{Y}\right)^{-1}\mathbf{u}$ [±$\begin{bmatrix}\pm\lambda\mathbf{I}_{Y}&\mathbf{H}\\ \mathbf{H}^{H}&\mp\mathbf{I}_{X}\end{bmatrix}$
@@
OMF $\lambda^{-1}\mathbf{H}^{H}\mathbf{u}$ [±$\begin{bmatrix}\pm\lambda\mathbf{I}_{Y}&\mathbf{0}_{Y\times X}\\ \mathbf{H}^{H}&\mp\mathbf{I}_{X}\end{bmatrix}$
@@ IV-A Arbitrary Beamforming
Consider a transmitter that linearly precodes a vector of $N_{S}$ symbols $\mathbf{s}\in\mathbb{C}^{N_{S}\times 1}$ to obtain the transmitted signal as $\mathbf{x}=\mathbf{W}\mathbf{s}$, where $\mathbf{W}\in\mathbb{C}^{N_{T}\times N_{S}}$ is an arbitrary precoding matrix. Such a transmitter can be implemented purely in the analog domain by a MiLAC having $N=N_{S}$ ports receiving the symbol vector in input and $M=N_{T}$ ports delivering the output to antennas perfectly matched to the reference impedance $Z_{0}=Y_{0}^{-1}$, as represented in Fig. 2. Since the antennas are perfectly matched, their Thevenin equivalent circuit is an impedance $Z_{0}=Y_{0}^{-1}$ connected to ground [36, Chapter 2.13]. Thus, the MiLAC in Fig. 2 behaves as when its last $N_{T}$ ports are connected to ground through an impedance $Z_{0}=Y_{0}^{-1}$, i.e., as analyzed in Section II. Following the analysis in Section II, such a MiLAC returns $\mathbf{x}=\mathbf{W}\mathbf{s}$ on its last $N_{T}$ ports when the signal $\mathbf{s}$ is fed as input on the first $N_{S}$ ports and the tunable admittance components are set such that the matrix $\mathbf{P}$ is
@@
for $k=1,\ldots,N_{S}+N_{T}$. From (19) and (20), we observe that the tunable admittance components of a MiLAC can be reconfigured in closed form to implement any arbitrary beamforming $\mathbf{W}$ matrix in the analog domain. The computational complexity of this beamforming technique, denoted as MiLAC-aided beamforming, is given by the complexity of designing $\mathbf{W}$ (depending on the adopted precoding strategy) and computing (19) and (20) (requiring in total $4N_{S}N_{T}$ real operations). Note that these operations are executed at each coherence block, while no operation is required on a per-symbol basis since the matrix-vector product $\mathbf{W}\mathbf{s}$ is performed in the analog domain.
@@
As discussed on the transmitter side, MiLAC-aided beamforming can also be introduced on the receiver side. Specifically, a MiLAC can be used to linearly combine the received signal $\mathbf{y}\in\mathbb{C}^{N_{R}\times 1}$ to obtain the signal used for detection $\mathbf{z}\in\mathbb{C}^{N_{S}\times 1}$ as $\mathbf{z}=\mathbf{G}\mathbf{y}$, where $\mathbf{G}\in\mathbb{C}^{N_{S}\times N_{R}}$ is an arbitrary combining matrix. This can be obtained with MiLAC having $N=N_{R}$ ports acquiring the input from the receiving antennas and $M=N_{S}$ ports providing the output vector $\mathbf{z}$, as represented in Fig. 3. Note that in Fig. 3 we assume the antennas to be perfectly matched to the reference impedance $Z_{0}=Y_{0}^{-1}$, such that their Thevenin equivalent circuit is a voltage generator (inducing a voltage $y_{n}$ at port $n$, for $n=1,\ldots,N$) in series to an impedance $Z_{0}=Y_{0}^{-1}$ [36, Chapter 2.13], i.e., as in the MiLAC analyzed in Section II.
@@ IV-B LMMSE-Inspired Beamforming at the Transmitter:
Assume that the previously considered MIMO system $\mathbf{y}=\mathbf{H}\mathbf{x}+\mathbf{n}$ is a multi-user system between an $N_{T}$-antenna transmitter and $N_{R}$ single-antenna users, with $N_{R}\leq N_{T}$. The signal received at the $N_{R}$ users $\mathbf{y}$ is used to detect the symbols $\mathbf{s}\in\mathbb{C}^{N_{S}\times 1}$, with $N_{S}=N_{R}$, and the transmitted signal is given by $\mathbf{x}=\mathbf{W}\mathbf{s}$, where $\mathbf{W}\in\mathbb{C}^{N_{T}\times N_{R}}$ is the precoding matrix implemented by a MiLAC, as in Fig. 2. By exploiting the operations included in Tab. III, it is possible to reconfigure the microwave network of the MiLAC to efficiently execute three popular linear transmitters [35, Chapter 12]. First, a MiLAC computing RLS can realize R-ZFBF, with precoding matrix $\mathbf{W}_{\text{R-ZFBF}}=\mathbf{H}^{H}(\mathbf{H}\mathbf{H}^{H}+\lambda\mathbf{I}_{N_{R}})^{-1}$, where $\lambda$ is an arbitrary regularization factor. Second, ZFBF can be implemented by a MiLAC computing OLS, whose precoding matrix is $\mathbf{W}_{\text{ZFBF}}=\mathbf{H}^{H}(\mathbf{H}\mathbf{H}^{H})^{-1}$. Third, a MiLAC computing OMF can realize MBF, with precoding matrix given by $\mathbf{W}_{\text{MBF}}=\lambda^{-1}\mathbf{H}^{H}$. These three transmitters are obtained by reconfiguring the tunable admittance components of the MiLAC according to the channel $\mathbf{H}$ on a per coherence block basis, and do not require any operation on a per-symbol basis.
@@ IV-C LMMSE-Inspired Beamforming at the Receiver:
Assume that the MIMO system $\mathbf{y}=\mathbf{H}\mathbf{x}+\mathbf{n}$ is a single-user system between an $N_{T}$-antenna transmitter and a $N_{R}$-antenna receiver, with $N_{R}\geq N_{T}$. The transmitted signal $\mathbf{x}$ contains the symbols to be detected at the receiver, i.e., $\mathbf{x}=\mathbf{s}$, and the receiver uses the signal $\mathbf{z}=\mathbf{G}\mathbf{y}\in\mathbb{C}^{N_{T}\times 1}$ to accurately detect the symbols, where $\mathbf{G}\in\mathbb{C}^{N_{T}\times N_{R}}$ is the combining matrix imposed by the MiLAC, as in Fig. 3. By using the operations in Tab. II, it is possible to reconfigure the microwave network of the MiLAC to efficiently implement three widely used receivers [35, Chapter 7]. First, a MiLAC computing RLS can realize the MMSE receiver with combining matrix $\mathbf{G}_{\text{MMSE}}=(\mathbf{H}^{H}\mathbf{H}+\lambda\mathbf{I}_{N_{T}})^{-1}\mathbf{H}^{H}$. If the covariance matrix of the transmitted signal $\mathbf{x}$ is $\mathbf{C}_{\mathbf{x}}=P_{T}/N_{T}\mathbf{I}_{N_{T}}$, where $P_{T}$ is the transmit power, and the covariance matrix of the noise $\mathbf{n}$ is $\mathbf{C}_{\mathbf{n}}=\sigma_{\mathbf{n}}^{2}\mathbf{I}_{N_{R}}$, where $\sigma_{\mathbf{n}}^{2}$ is the noise power, the optimal regularization factor $\lambda$ is given by $\lambda=N_{T}\sigma_{\mathbf{n}}^{2}/P_{T}$ [35, Chapter 7]. Second, the ZF receiver can be realized by a MiLAC computing OLS, whose combining matrix is $\mathbf{G}_{\text{ZF}}=(\mathbf{H}^{H}\mathbf{H})^{-1}\mathbf{H}^{H}$. Third, a MiLAC computing OMF can implement the MF receiver, with combining matrix $\mathbf{G}_{\text{MF}}=\lambda^{-1}\mathbf{H}^{H}$, where the optimal $\lambda$ is $\lambda=N_{T}\sigma_{\mathbf{n}}^{2}/P_{T}$ if $\mathbf{C}_{\mathbf{x}}=P_{T}/N_{T}\mathbf{I}_{N_{T}}$ and $\mathbf{C}_{\mathbf{n}}=\sigma_{\mathbf{n}}^{2}\mathbf{I}_{N_{R}}$ [35, Chapter 7]. These three receivers are realized by solely reconfiguring the MiLAC on a per coherence block basis, while no operation is required on a per-symbol basis.
@@
IV-D Computing the DFT
for $i\neq k$, and
$Y_{k,k}=\begin{cases}-Y_{0}\sqrt{N_{R}}&k=1\\ 0&k\neq 1\end{cases},$ (24)
for $k=1,\ldots,2N_{R}$.
Remark 4.

### v3

@@ III-C Computing the Special Cases of the LMMSE Estimator with a MiLAC
The expressions of the five special cases of the LMMSE estimator can be efficiently returned on the output vector $\mathbf{v}_{2}$ of a MiLAC having $N=Y$ input ports and $M=X$ output ports with no input, i.e., $P=X+Y$ total ports. This is achieved by exploiting (5) and (7), as discussed for the LMMSE estimator in Part I. First, the tunable admittance components $\{Y_{i,k}\}_{i,k=1}^{P}$ are set according to (8) such that the blocks of the matrix $\mathbf{P}$, namely $\mathbf{P}_{11}$, $\mathbf{P}_{12}$, $\mathbf{P}_{21}$, and $\mathbf{P}_{22}$, give the desired expression in the output $\mathbf{v}_{2}$. Second, the input vector of the MiLAC is set as $\mathbf{u}=\mathbf{y}$. In detail, the expressions of the special cases in (11)-(18) can be computed by setting the blocks of the matrix $\mathbf{P}$ as illustrated in Tabs. II and III, depending on the invertibility of the blocks $\mathbf{P}_{11}$ and $\mathbf{P}_{22}$. When the block $\mathbf{P}_{11}$ is invertible, $\mathbf{v}_{2}$ is given by (5), as reported in Tab. II, while when the block $\mathbf{P}_{22}$ is invertible, $\mathbf{v}_{2}$ is given by (7), as reported in Tab. III. Note that GLS and OLS are computed via a matrix $\mathbf{P}$ whose blocks $\mathbf{P}_{11}$ and $\mathbf{P}_{22}$ are not both invertible. Thus, for GLS and OLS the expressions in Tabs. II and III are not equivalent in general, while they are equivalent for the LMMSE estimator and the other special cases.
@@
The five special cases of the LMMSE estimator can be efficiently computed using a MiLAC, which offers significantly reduced computational complexity compared to a digital computer, in terms of the number of arithmetic real operations required. This is because a MiLAC directly performs matrix-matrix multiplications and matrix inversion operations in the analog domain. On the one hand, the computational complexity of computing the special cases with a MiLAC is driven by the number of operations required to compute (8) for a given matrix $\mathbf{P}$. Thus, the complexity required to compute GLS, RLS, and OLS with a MiLAC is the same as the complexity of computing the LMMSE estimator, i.e., $6XY$ real operations, as discussed in Part I. Besides, computing (8) for the GMF and OMF require approximately $2XY$ real operations for both cases $i\neq k$ and $i=k$, resulting in a total of $4XY$ real operations, where we assumed that the operations involving $\mathbf{C}_{\mathbf{x}}$ and $\mathbf{C}_{\mathbf{n}}$ are precomputed offline. On the other hand, a digital computer calculates the GLS with the same complexity as the LMMSE estimator due to the two matrix-matrix products and matrix inversion, i.e., $8(XY^{2}+X^{2}Y+\min\{X^{3},Y^{3}\}/3)$, as discussed in Part I. The RLS and OLS can be computed with reduced complexity as they require only one matrix-matrix product and a matrix inversion, namely $8\min\{X^{2}Y+X^{3}/3,XY^{2}+Y^{3}/3\}$. Besides, the GMF and OMF only require three and one matrix-vector products, respectively, resulting in a complexity of $8(X^{2}+XY+Y^{2})$ and $8XY$ real operations, respectively.
@@
LMMSE $\mathbf{C}_{\mathbf{x}}\mathbf{H}^{H}\left(\mathbf{H}\mathbf{C}_{\mathbf{x}}\mathbf{H}^{H}+\mathbf{C}_{\mathbf{n}}\right)^{-1}\mathbf{u}$ $\begin{bmatrix}\pm\mathbf{C}_{\mathbf{n}}&\mathbf{H}\\ \mathbf{H}^{H}&\mp\mathbf{C}_{\mathbf{x}}^{-1}\end{bmatrix}$
@@
GMF $\mathbf{C}_{\mathbf{x}}\mathbf{H}^{H}\mathbf{C}_{\mathbf{n}}^{-1}\mathbf{u}$ $\begin{bmatrix}\pm\mathbf{C}_{\mathbf{n}}&\mathbf{0}_{Y\times X}\\ \mathbf{H}^{H}&\mp\mathbf{C}_{\mathbf{x}}^{-1}\end{bmatrix}$
RLS $\mathbf{H}^{H}\left(\mathbf{H}\mathbf{H}^{H}+\lambda\mathbf{I}_{Y}\right)^{-1}\mathbf{u}$ $\begin{bmatrix}\pm\lambda\mathbf{I}_{Y}&\mathbf{H}\\ \mathbf{H}^{H}&\mp\mathbf{I}_{X}\end{bmatrix}$
@@
OMF $\lambda^{-1}\mathbf{H}^{H}\mathbf{u}$ $\begin{bmatrix}\pm\lambda\mathbf{I}_{Y}&\mathbf{0}_{Y\times X}\\ \mathbf{H}^{H}&\mp\mathbf{I}_{X}\end{bmatrix}$
@@ IV-A Arbitrary Beamforming
Consider a transmitter that linearly precodes a vector of $N_{S}$ symbols $\mathbf{s}\in\mathbb{C}^{N_{S}\times 1}$ to obtain the transmitted signal as $\mathbf{x}=\mathbf{W}\mathbf{s}$, where $\mathbf{W}\in\mathbb{C}^{N_{T}\times N_{S}}$ is an arbitrary precoding matrix. Such a transmitter can be implemented purely in the analog domain by a MiLAC having $N=N_{S}$ ports receiving the symbol vector in input and $M=N_{T}$ ports delivering the output to antennas perfectly matched to the reference impedance $Z_{0}=Y_{0}^{-1}$, as represented in Fig. 2. Since the antennas are perfectly matched, their Thevenin equivalent circuit is an impedance $Z_{0}=Y_{0}^{-1}$ connected to ground [45, Chapter 2.13]. Thus, the MiLAC in Fig. 2 behaves as when its last $N_{T}$ ports are connected to ground through an impedance $Z_{0}=Y_{0}^{-1}$, i.e., as analyzed in Section II. Following the analysis in Section II, such a MiLAC returns $\mathbf{x}=\mathbf{W}\mathbf{s}$ on its last $N_{T}$ ports when the signal $\mathbf{s}$ is fed as input on the first $N_{S}$ ports and the tunable admittance components are set such that the matrix $\mathbf{P}$ is
@@
for $k=1,\ldots,N_{S}+N_{T}$. From (20) and (21), we observe that the tunable admittance components of a MiLAC can be reconfigured in closed form to implement any arbitrary beamforming $\mathbf{W}$ matrix in the analog domain. The computational complexity of this beamforming technique, denoted as MiLAC-aided beamforming, is given by the complexity of designing $\mathbf{W}$ (depending on the adopted precoding strategy) and computing (20) and (21) (requiring in total $4N_{S}N_{T}$ real operations). Note that these operations are executed at each coherence block, while no operation is required on a per-symbol basis since the matrix-vector product $\mathbf{W}\mathbf{s}$ is performed in the analog domain. To adapt the beamforming matrix $\mathbf{W}$ to the wireless channel $\mathbf{H}$, the MiLAC is assumed to be implemented with tunable admittance components that can be reconfigured at every channel coherence time.
@@
As discussed on the transmitter side, MiLAC-aided beamforming can also be introduced on the receiver side. Specifically, a MiLAC can be used to linearly combine the received signal $\mathbf{y}\in\mathbb{C}^{N_{R}\times 1}$ to obtain the signal used for detection $\mathbf{z}\in\mathbb{C}^{N_{S}\times 1}$ as $\mathbf{z}=\mathbf{G}\mathbf{y}$, where $\mathbf{G}\in\mathbb{C}^{N_{S}\times N_{R}}$ is an arbitrary combining matrix. This can be obtained with MiLAC having $N=N_{R}$ ports acquiring the input from the receiving antennas and $M=N_{S}$ ports providing the output vector $\mathbf{z}$, as represented in Fig. 3. Note that in Fig. 3 we assume the antennas to be perfectly matched to the reference impedance $Z_{0}=Y_{0}^{-1}$, such that their Thevenin equivalent circuit is a voltage generator (inducing a voltage $y_{n}$ at port $n$, for $n=1,\ldots,N$) in series to an impedance $Z_{0}=Y_{0}^{-1}$ [45, Chapter 2.13], i.e., as in the MiLAC analyzed in Section II.
@@ IV-B LMMSE-Inspired Beamforming at the Transmitter:
Assume that the previously considered MIMO system $\mathbf{y}=\mathbf{H}\mathbf{x}+\mathbf{n}$ is a multi-user system between an $N_{T}$-antenna transmitter and $N_{R}$ single-antenna users, with $N_{R}\leq N_{T}$. The signal received at the $N_{R}$ users $\mathbf{y}$ is used to detect the symbols $\mathbf{s}\in\mathbb{C}^{N_{S}\times 1}$, with $N_{S}=N_{R}$, and the transmitted signal is given by $\mathbf{x}=\mathbf{W}\mathbf{s}$, where $\mathbf{W}\in\mathbb{C}^{N_{T}\times N_{R}}$ is the precoding matrix implemented by a MiLAC, as in Fig. 2. By exploiting the operations included in Tab. III, it is possible to reconfigure the microwave network of the MiLAC to efficiently execute three popular linear transmitters [44, Chapter 12]. First, a MiLAC computing RLS can realize R-ZFBF, with precoding matrix $\mathbf{W}_{\text{R-ZFBF}}=\mathbf{H}^{H}(\mathbf{H}\mathbf{H}^{H}+\lambda\mathbf{I}_{N_{R}})^{-1}$, where $\lambda$ is an arbitrary regularization factor. Second, ZFBF can be implemented by a MiLAC computing OLS, whose precoding matrix is $\mathbf{W}_{\text{ZFBF}}=\mathbf{H}^{H}(\mathbf{H}\mathbf{H}^{H})^{-1}$. Third, a MiLAC computing OMF can realize MBF, with precoding matrix given by $\mathbf{W}_{\text{MBF}}=\lambda^{-1}\mathbf{H}^{H}$. These three transmitters are obtained by reconfiguring the tunable admittance components of the MiLAC according to the channel $\mathbf{H}$ on a per coherence block basis, and do not require any operation on a per-symbol basis. In addition, they require ultra-low computational complexity on a per coherence block basis since the matrix $\mathbf{P}$ of the MiLAC is readily given as a function of the wireless channel $\mathbf{H}$. Hence, the only operations required are to compute the admittance values $\{Y_{i,k}\}_{i,k=1}^{P}$ from $\mathbf{P}$.
@@ IV-C LMMSE-Inspired Beamforming at the Receiver:
Assume that the MIMO system $\mathbf{y}=\mathbf{H}\mathbf{x}+\mathbf{n}$ is a single-user system between an $N_{T}$-antenna transmitter and a $N_{R}$-antenna receiver, with $N_{R}\geq N_{T}$. The transmitted signal $\mathbf{x}$ contains the symbols to be detected at the receiver, i.e., $\mathbf{x}=\mathbf{s}$, and the receiver uses the signal $\mathbf{z}=\mathbf{G}\mathbf{y}\in\mathbb{C}^{N_{T}\times 1}$ to accurately detect the symbols, where $\mathbf{G}\in\mathbb{C}^{N_{T}\times N_{R}}$ is the combining matrix imposed by the MiLAC, as in Fig. 3. By using the operations in Tab. II, it is possible to reconfigure the microwave network of the MiLAC to efficiently implement three widely used receivers [44, Chapter 7]. First, a MiLAC computing RLS can realize the MMSE receiver with combining matrix $\mathbf{G}_{\text{MMSE}}=(\mathbf{H}^{H}\mathbf{H}+\lambda\mathbf{I}_{N_{T}})^{-1}\mathbf{H}^{H}$. If the covariance matrix of the transmitted signal $\mathbf{x}$ is $\mathbf{C}_{\mathbf{x}}=P_{T}/N_{T}\mathbf{I}_{N_{T}}$, where $P_{T}$ is the transmit power, and the covariance matrix of the noise $\mathbf{n}$ is $\mathbf{C}_{\mathbf{n}}=\sigma_{\mathbf{n}}^{2}\mathbf{I}_{N_{R}}$, where $\sigma_{\mathbf{n}}^{2}$ is the noise power, the optimal regularization factor $\lambda$ is given by $\lambda=N_{T}\sigma_{\mathbf{n}}^{2}/P_{T}$ [44, Chapter 7]. Second, the ZF receiver can be realized by a MiLAC computing OLS, whose combining matrix is $\mathbf{G}_{\text{ZF}}=(\mathbf{H}^{H}\mathbf{H})^{-1}\mathbf{H}^{H}$. Third, a MiLAC computing OMF can implement the MF receiver, with combining matrix $\mathbf{G}_{\text{MF}}=\lambda^{-1}\mathbf{H}^{H}$, where the optimal $\lambda$ is $\lambda=N_{T}\sigma_{\mathbf{n}}^{2}/P_{T}$ if $\mathbf{C}_{\mathbf{x}}=P_{T}/N_{T}\mathbf{I}_{N_{T}}$ and $\mathbf{C}_{\mathbf{n}}=\sigma_{\mathbf{n}}^{2}\mathbf{I}_{N_{R}}$ [44, Chapter 7]. These three receivers are realized by solely reconfiguring the MiLAC on a per coherence block basis, while no operation is required on a per-symbol basis.
@@
IV-D Computing the DFT
for $i\neq k$, and
$Y_{k,k}=\begin{cases}-Y_{0}\sqrt{N_{R}}&k=1\\ 0&k\neq 1\end{cases},$ (25)
for $k=1,\ldots,2N_{R}$.
Remark 4.